OpenStudy (anonymous):

How many zeros does 6z^7+iz have and how many are real?

OpenStudy (jamesj):

Well 6z^7+iz = 0 if and only if z (z^6 - (-i/6)) = 0 Hence there's one clear real zero, z = 0 But now, how many complex roots of z^6 = -i/6 are real?

OpenStudy (anonymous):

what

OpenStudy (jamesj):

Do you agree with this first: 6z^7+iz = 0 if and only if z (z^6 - (-i/6)) = 0

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Now z (z^6 - (-i/6)) = 0 iff z = 0 or z^6 - (-i/6) = 0

OpenStudy (jamesj):

Hence z = 0 is a zero, which makes sense. Now there are six other roots from the equation z^6 - (-i/6) = 0

OpenStudy (anonymous):

ok